# Improved Search-to-Decision Reduction for Random Local Functions

### v1

@@ 3 Search-to-Decision Reduction
Suppose that $P$ is $c$-correlated, we can then assume that $m=o(n^{c/2-2})$ as if not then both Search and Decision for distribution $G_{n,k\times m,d}$ will become easy due to Lemma 2.13, which our theorem statement is then trivially true. The same reasoning implies that we can assume $\varepsilon=w(n^{-c/4})$. This assumption will be used in Claim 3.10
@@ 3.1 Hybrids
Suppose there are $t$ distributions $H_{0},\dots,H_{t}$ such that an algorithm $\mathsf{D}$ has advatange $\varepsilon$ in distinguishing $H_{0}$ and $H_{t}$. Then the algorithm can distinguish $H_{i}$ and $H_{i+1}$ with advantage $\varepsilon/t$ on a randomly chosen $i$. Formally,
@@
Define a randomized transformation $T_{a,b}:G_{n,m,d}\to G_{n,m,d}$ as follows: On a input a hypergraph $G$, each of the hyperedges $S_{i}=(j_{1},\dots,j_{d})$ is transformed to $S_{i}^{\prime}=(j_{1}^{\prime},\dots,j_{d}^{\prime})$ independently where
@@
For $H_{0}^{s}$, the only changes made are that a randomly sampled permutation is applied to the hypergraph. However, since the hypergraph is randomly sampled and $s^{\prime}$ is sampled such that it has the same hamming weight as $s$, it is identically distributed as $(G,f_{G,P}(s^{\prime}))$.
@@ 3.2 Predictor
Consider any $d$-ary predicate $P$ with bounded bias $\eta$, and suppose there is a polynomial-time algorithm $\mathsf{D}$ that has advantage $\varepsilon$ to solve the Decision Problem for $(P,n,m)$. Then there is a set of polynomial-time algorithms $\{\mathsf{S}_{2},\dots,\mathsf{S}_{n}\}$ such that for all $\varepsilon$-fairly balanced and good $s\in\mathcal{G}_{D}$, there are a pair of numbers $eq_{s},neq_{s}\in[0,1]$ and the algorithms behave as follows for each $i\in[2,n]$:
@@
  1. 1.
    Sample a random number $r\leftarrow[0,t-1]$.
  2. 2.
    Sample a random permutation $\pi$ on $[n]$.
  3. 3.
    Get $H_{r}^{s}$ by performing $r$ hybridization steps on $G$, i.e. $H_{r}^{s}=(T_{a_{r},b_{r}}\circ\ldots\circ T_{a_{1},b_{1}}(\pi(G)),f_{G,P}(s))$, where $a_{j},b_{j}\leftarrow[n]$.
  4. 4.
    Apply $T_{\pi(1),\pi(i)}$ to $H_{r}^{s}$ to obtain $H$.
  5. 5.
    Return $\mathsf{D}(H)$.
@@
When $s_{1}=s_{i}$, this implies $\pi(s)_{\pi(1)}=\pi(s)_{\pi(i)}$, which means that the transformation $T_{\pi(1),\pi(i)}$ on the hypergraph does not affect the output of the $d$-local function. Essentially, no effective randomisation is performed, and the distribution does not look more like the null distribution. Therefore we are be able to conclude that $H\approx H_{r}^{s}$. Details are deferred to Section 5.4
@@ 3.3 Amplification
The estimation can be done by locally generating a secret of a given hamming weight and a random problem instance, then applying hybridisation and sending it to the distinguisher $\mathsf{D}$. With $O((t/\varepsilon)^{4}\log(n/\varepsilon))$ number of trials, by Hoeffding’s inequality, the average will converge to within $o(\varepsilon/t)$ of the true value of $eq_{s}$, with failure probability $o(\varepsilon)$. Once we have these estimate, we run the computations below with each of them until the correct one is used and the secret $s$ is found.
@@
Now that we have a predictor of $s_{i}$ with high probability, use the same set of problem instances to learn other $s_{j}$ for $j\in[2,n]$. Perform a union bound over all $i$, we set $l=\Theta(t^{2}\log(n/\varepsilon)/\varepsilon^{2})$ for a failure probability of at most $\varepsilon/4$ in guessing correctly whether $s_{i}=s_{1}$ for all $i\in[2,n]$.
Finally, we can now prove our main theorem Theorem 3.1. Since we assumed that $d=O(1)$, the $d$-ary $c$-correlated predicate $P$ must have bounded bias (depending on $d$), and we can assume $m=o(n^{c/2-2})$ and $\varepsilon=w(n^{-c/4})$. The recovery of the secret in total will cost $O((n^{2}/\varepsilon^{2})\log^{3}(n/\varepsilon))\times m$ samples and runs in time polynomial of $n,m$ and $T$ (runtime of decision algorithm). The success probability is determined by the number of good secrets and whether the secret is $\varepsilon$-fairly balanced, as the above algorithm works if both these conditions are met. The probability that a randomly selected secret is not good is $1-\varepsilon/2$ (Claim 3.13), the probability that a randomly selected secret is not fairly balanced is $o(\varepsilon)$ (Remark 3.7). By union bound, the probability that a randomly selected secret is both good and fairly balanced and the algorithm succeeds is $\Omega(\varepsilon)$. Finally, if the original distinguisher runs in polynomial time, then all of the above can be done in polynomial time as well. This proves Theorem 3.1.
@@ 4.1 Non-Constant Sparsity
The second change is in the final part where we show the statistical distance between $f_{G,P}(s)$ and $Bern(\eta)^{m}$ is $o(\varepsilon)$, where we will use the asumption that $(m/n^{c})\cdot(\log^{r}(n)\log(1/\varepsilon))^{2c}=o(\varepsilon^{2})$ as follows.
@@ 5.1 Random Graph
This is because $T_{a,b}$ only changes the value of a vertex in a hypergraph if it is either $a$ or $b$. Otherwise, it remains unchanged. Our goal is to prove that the random averaging process converges to a uniform distribution; namely, when $t=\Omega(n\log(n/\varepsilon))$, $\sum_{k\in[n]}(\left|x_{k}(t)-1/n\right|)$ becomes small. This then somewhat resembles the randomised gossip process, which we can then use a similar approach from [BGPS06] to show convergence. To prove that, we use the $L_{2}$ deviation function $V$ defined as
@@
Each vertices in the hyperedge has their own distribution $D$. Note that their distributions for different vertices are not independent due to the choice of $a_{j},b_{j}$. It could also not be identitical due to different initial states. Nevertheless, each vertex is sampled independently from its own distribution $D$.
We say that a distribution of a vertex is bad if the $(a_{j},b_{j})$’s are chosen in such a way that $V(t)$ for that vertex is larger than $z=(md/\varepsilon^{2})\exp(-t/n)$. The probability that the $D$ is bad is at most $\varepsilon^{2}/md$ (Eq. 2). By union bound, across the $md$ vertices in all the hyperedges, there is a distribution that is bad with probability at most $\varepsilon^{2}$. In the case of all distributions being good, the distance from uniform for one the vertices is at most $\sqrt{nz}$. Therefore, the overall statistical distance between the product distribution and a randomly sampled hypergraph can be upperbounded by $\varepsilon^{2}+md\cdot\sqrt{nz}$
@@ 5.2 Random Output
Since $s$ is $\varepsilon$-fairly balance, $s$ has hamming weight bounded by [n/2w,n/2+w][n/2-w,n/2+w] where $w=2\sqrt{n}\log(1/\varepsilon)$. Without loss of generality, assume that $s$ has at least as many ones as zeros. Let $D$ be the distribution of randomly sampling an element in the binary string $s$. Then, $D$ is a Bernoulli distribution that is only slightly biased.
@@ 5.3 Good Secret
Therefore there should be at least $\varepsilon/2$ fractions of secret $s\in\mathbb{F}_{2}^{n}$ such that $s\in\mathcal{G}_{\mathsf{D}}$, which implies $|\mathcal{G}_{\mathsf{D}}|\geq(\varepsilon/2)\cdot 2^{n}$
@@ 5.4 Equal
Define a deterministic transformation : $T^{\prime}_{a,b}:\mathbb{F}_{2}^{md}\times G_{n,m,d}\to G_{n,m,d}$. On a input of a vector $v\in\mathbb{F}_{2}^{md}$ and a hypergraph $G$, each of the hyperedges $S_{i}=(j_{1},\dots,j_{d})$ is transformed to $S_{i}^{\prime}=(j_{1}^{\prime},\dots,j_{d}^{\prime})$ as follows:
@@ 5.5 Not Equal
This is because in the last transformation that we apply to $H_{r}^{s}$ to get $H$, our assumption on $s_{1}\neq s_{i}$ ensures that $T_{\pi(1),\pi(i)}$ always has $s^{\prime}_{\pi(1)}\neq s^{\prime}_{\pi(i)}$. Futhermore, the randomness of $\pi$ ensures that $(\pi(1),\pi(i))$ is distributed just as sampling $(a,b)$ condition on $s^{\prime}_{a}\neq s^{\prime}_{b}$. Using Eq. 3,
@@ Acknowledgements
Both authors are supported by the National Research Foundation, Singapore, under its NRF Fellowship programme, award no. NRF-NRFF14-2022-0010. AI tools were used as typing assistants for grammar and basic editing.

### v2

@@ 3 Search-to-Decision Reduction
Suppose that $P$ is $c$-correlated, we can then assume that $m=o(n^{c/2-2})$ as if not then both Search and Decision for distribution $G_{n,k\times m,d}$ will become easy due to Lemma 2.13, making our theorem statement trivially true. The same reasoning implies that we can assume $\varepsilon=w(n^{-c/4})$. This assumption will be used in Claim 3.10
@@ 3.1 Hybrids
Suppose there are $t$ distributions $H_{0},\dots,H_{t}$ such that an algorithm $\mathsf{D}$ has advantage $\varepsilon$ in distinguishing $H_{0}$ and $H_{t}$. Then the algorithm can distinguish $H_{i}$ and $H_{i+1}$ with advantage $\varepsilon/t$ on a randomly chosen $i$. Formally,
@@
Define a randomized transformation $T_{a,b}:G_{n,m,d}\to G_{n,m,d}$ as follows: On inputting a hypergraph $G$, each of the hyperedges $S_{i}=(j_{1},\dots,j_{d})$ is transformed to $S_{i}^{\prime}=(j_{1}^{\prime},\dots,j_{d}^{\prime})$ independently where
@@
For $H_{0}^{s}$, the only changes made are that a randomly sampled permutation is applied to the hypergraph. However, since the hypergraph is randomly sampled and $s^{\prime}$ is sampled such that it has the same Hamming weight as $s$, it is identically distributed to $(G,f_{G,P}(s^{\prime}))$.
@@ 3.2 Predictor
Consider any $d$-ary predicate $P$ with bounded bias $\eta$, and suppose there is a polynomial-time algorithm $\mathsf{D}$ that has advantage $\varepsilon$ to solve the Decision Problem for $(P,n,m)$. Then there is a set of polynomial-time algorithms $\{\mathsf{S}_{2},\dots,\mathsf{S}_{n}\}$ such that for all $\varepsilon$-fairly balanced and good $s\in\mathcal{G}_{D}$, there is a pair of numbers $eq_{s},neq_{s}\in[0,1]$ and the algorithms behave as follows for each $i\in[2,n]$:
@@
Algorithm $\mathsf{S}_{i}(G,f_{G,P}(s))$: 1. Sample a random number $r\leftarrow[0,t-1]$. 2. Sample a random permutation $\pi$ on $[n]$. 3. Get $H_{r}^{s}$ by performing $r$ hybridization steps on $G$, i.e.
$H_{r}^{s}=(T_{a_{r},b_{r}}\circ\ldots\circ T_{a_{1},b_{1}}(\pi(G)),f_{G,P}(s))$, where $a_{j},b_{j}\leftarrow[n]$.
4. Apply $T_{\pi(1),\pi(i)}$ to $H_{r}^{s}$ to obtain $H$. 5. Return $\mathsf{D}(H)$.
@@
When $s_{1}=s_{i}$, this implies $\pi(s)_{\pi(1)}=\pi(s)_{\pi(i)}$, which means that the transformation $T_{\pi(1),\pi(i)}$ on the hypergraph does not affect the output of the $d$-local function. Essentially, no effective randomisation is performed, and the distribution does not look more like the null distribution. Therefore we are able to conclude that $H\approx H_{r}^{s}$. Details are deferred to Section 5.4
@@ 3.3 Amplification
The estimation can be done by locally generating a secret of a given Hamming weight and a random problem instance, then applying hybridisation and sending it to the distinguisher $\mathsf{D}$. With $O((t/\varepsilon)^{4}\log(n/\varepsilon))$ number of trials, by Hoeffding’s inequality, the average will converge to within $o(\varepsilon/t)$ of the true value of $eq_{s}$, with failure probability $o(\varepsilon)$. Once we have these estimates, we run the computations below with each of them until the correct one is used and the secret $s$ is found.
@@
Now that we have a predictor of $s_{i}$ with high probability, use the same set of problem instances to learn other $s_{j}$ for $j\in[2,n]$. Performing a union bound over all $i$, we set $l=\Theta(t^{2}\log(n/\varepsilon)/\varepsilon^{2})$ for a failure probability of at most $\varepsilon/4$ in guessing correctly whether $s_{i}=s_{1}$ for all $i\in[2,n]$.
Finally, we can now prove our main theorem Theorem 3.1. Since we assumed that $d=O(1)$, the $d$-ary $c$-correlated predicate $P$ must have bounded bias (depending on $d$), and we can assume $m=o(n^{c/2-2})$ and $\varepsilon=w(n^{-c/4})$. The recovery of the secret in total costs $O((n^{2}/\varepsilon^{2})\log^{3}(n/\varepsilon))\times m$ samples and runs in time polynomial of $n,m$ and $T$ (runtime of decision algorithm). The success probability is determined by the number of good secrets and whether the secret is $\varepsilon$-fairly balanced, as the above algorithm works if both these conditions are met. The probability that a randomly selected secret is not good is $1-\varepsilon/2$ (Claim 3.13), the probability that a randomly selected secret is not fairly balanced is $o(\varepsilon)$ (Remark 3.7). By union bound, the probability that a randomly selected secret is both good and fairly balanced and the algorithm succeeds is $\Omega(\varepsilon)$. Finally, if the original distinguisher runs in polynomial time, then all of the above can be done in polynomial time as well. This proves Theorem 3.1.
@@ 4.1 Non-Constant Sparsity
The second change is in the final part where we show the statistical distance between $f_{G,P}(s)$ and $Bern(\eta)^{m}$ is $o(\varepsilon)$, where we will use the assumption that $(m/n^{c})\cdot(\log^{r}(n)\log(1/\varepsilon))^{2c}=o(\varepsilon^{2})$ as follows.
@@ 5.1 Random Graph
This is because $T_{a,b}$ only changes the value of a vertex in a hypergraph if it is either $a$ or $b$. Otherwise, it remains unchanged. Our goal is to prove that the random averaging process converges to a uniform distribution; namely, when $t=\Omega(n\log(n/\varepsilon))$, $\sum_{k\in[n]}(\left|x_{k}(t)-1/n\right|)$ becomes small. This then somewhat resembles the randomised gossip process, which we can then use a similar approach from [BGP+06] to show convergence. To prove that, we use the $L_{2}$ deviation function $V$ defined as
@@
Each vertex in the hyperedge has its own distribution $D$. Note that their distributions for different vertices are not independent due to the choice of $a_{j},b_{j}$. It could also not be identical due to different initial states. Nevertheless, each vertex is sampled independently from its own distribution $D$.
We say that a distribution of a vertex is bad if the $(a_{j},b_{j})$’s are chosen in such a way that $V(t)$ for that vertex is larger than $z=(md/\varepsilon^{2})\exp(-t/n)$. The probability that the $D$ is bad is at most $\varepsilon^{2}/md$ (Eq. 2). By union bound, across the $md$ vertices in all the hyperedges, there is a distribution that is bad with probability at most $\varepsilon^{2}$. In the case of all distributions being good, the distance from uniform for one of the vertices is at most $\sqrt{nz}$. Therefore, the overall statistical distance between the product distribution and a randomly sampled hypergraph can be upperbounded by $\varepsilon^{2}+md\cdot\sqrt{nz}$
@@ 5.2 Random Output
Since $s$ is $\varepsilon$-fairly balanced, $s$ has Hamming weight bounded by [n/2w,n/2+w][n/2-w,n/2+w] where $w=2\sqrt{n}\log(1/\varepsilon)$. Without loss of generality, assume that $s$ has at least as many ones as zeros. Let $D$ be the distribution of randomly sampling an element in the binary string $s$. Then, $D$ is a Bernoulli distribution that is only slightly biased.
@@ 5.3 Good Secret
Therefore there should be at least $\varepsilon/2$ fraction of secret $s\in\mathbb{F}_{2}^{n}$ such that $s\in\mathcal{G}_{\mathsf{D}}$, which implies $|\mathcal{G}_{\mathsf{D}}|\geq(\varepsilon/2)\cdot 2^{n}$
@@ 5.4 Equal
Define a deterministic transformation : $T^{\prime}_{a,b}:\mathbb{F}_{2}^{md}\times G_{n,m,d}\to G_{n,m,d}$. On input of a vector $v\in\mathbb{F}_{2}^{md}$ and a hypergraph $G$, each of the hyperedges $S_{i}=(j_{1},\dots,j_{d})$ is transformed to $S_{i}^{\prime}=(j_{1}^{\prime},\dots,j_{d}^{\prime})$ as follows:
@@ 5.5 Not Equal
This is because in the last transformation that we apply to $H_{r}^{s}$ to get $H$, our assumption on $s_{1}\neq s_{i}$ ensures that $T_{\pi(1),\pi(i)}$ always has $s^{\prime}_{\pi(1)}\neq s^{\prime}_{\pi(i)}$. Furthermore, the randomness of $\pi$ ensures that $(\pi(1),\pi(i))$ is distributed just as sampling $(a,b)$ condition on $s^{\prime}_{a}\neq s^{\prime}_{b}$. Using Eq. 3,
@@ Acknowledgements
Both authors are supported by the National Research Foundation, Singapore, under its NRF Fellowship programme, award no. NRF-NRFF14-2022-0010. AI tools were used as typing assistants for grammar and basic editing. We thank the EUROCRYPT 2026 reviewers for their helpful suggestions.